OpenStudy (akohl):

@Abhisar A block of mass m = 4.4 kg slides from left to right across a frictionless surface with a speed 8.4 m/s It collides in a perfectly elastic collision with a second block of mass M that is at rest. After the collision, the 4.4-kg block reverses direction, and its new speed is 2.5 m/s What is V, the speed of the second block after the collision? 6.4 m/s 5.1 m/s 5.9 m/s 7.2 m/s

OpenStudy (abhisar):

Ok..let's see !

OpenStudy (abhisar):

Have u tried it yourself ?

OpenStudy (akohl):

Ye I get stuck because there no mass for M so you can't get velocity

OpenStudy (akohl):

Any ideas?

OpenStudy (abhisar):

one min !

OpenStudy (akohl):

Starting velocity of other block is rest so 0

OpenStudy (akohl):

K

OpenStudy (abhisar):

So u tried to equate the initial kinetic energy with final kinetic energy. Can u write down the equations ?

OpenStudy (akohl):

I did m*v+m*v=m+m*v where its each mass of the block times its own velocity

OpenStudy (abhisar):

it should be 4.4*8.4 = 4.4*-2.5 + MV

OpenStudy (abhisar):

Now i can see the problem. There is no M given.

OpenStudy (akohl):

Exactly

OpenStudy (akohl):

Its elastic does that help

OpenStudy (abhisar):

Actually we can calculate it, but first we will have to derive the equation for velocities by using conservation of kinetic energy and conservation of momentum

OpenStudy (akohl):

How

OpenStudy (abhisar):

Let's denote the initial velocity by u, final velocities by v1 and v2.

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar Now for conservation of momentum, \(\huge\tt mu = mv_1 + Mv_2\) --> \(\huge\tt m(u-v_1) = Mv_2\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (akohl):

Won't we still get that we cant solve since theres no mass or velocity

OpenStudy (abhisar):

wait.....it's not done yet. Just tell me if u got the above modification ?

OpenStudy (akohl):

Yea but its still just 48=mv2

OpenStudy (abhisar):

Now for conservation of kinetic energy \(\huge\tt \frac{1}{2}mu^2 = \frac{1}{2}mv_1^2 + \frac{1}{2}mv_2^2\) -->\(\huge\tt m(u-v)(u+v) = mv_2^2\) by using \(\huge\tt (a-b)^2 = (a-b)(a+b)\)

OpenStudy (abhisar):

Getting ?

OpenStudy (akohl):

What is a and b

OpenStudy (abhisar):

The mathematical identity !

OpenStudy (abhisar):

Like (a+b)^2 = a^2 + b^2 + 2ab

OpenStudy (akohl):

Im gettong 283=mv and still cant solve

OpenStudy (abhisar):

There is also an identity (a-b)^2 = (a-b)(a+b)

OpenStudy (abhisar):

Wait !!!!

OpenStudy (abhisar):

We havn't derived the formula yet..you want me to directly give you the formula !

OpenStudy (akohl):

Plz

OpenStudy (abhisar):

LOL !

OpenStudy (akohl):

Lol

OpenStudy (abhisar):

LOL wait !

OpenStudy (abhisar):

You still need to know mass M

OpenStudy (akohl):

Dang

OpenStudy (abhisar):

is ur question seriously complete ?

OpenStudy (akohl):

Yea this is all the info

OpenStudy (abhisar):

uh-oh !

OpenStudy (abhisar):

@ganeshie8 may help you !